OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x-7/x+3 and g(x)=-3x-7/x-1

OpenStudy (anonymous):

\[f(x)=\frac{x-7}{x+3}\] replace \(x\) by \(\frac{-3-7}{x-1}\) and get \[f(g(x))=\frac{\frac{-3-7}{x-1}-7}{\frac{-3-7}{x-1}+3}\]

OpenStudy (anonymous):

bunch of algebra follows multiply top and bottom by \(x-1\) to clear the fraction then a raft of cancellation will ensue, ending only with \(x\)

OpenStudy (anonymous):

can you show me its very confusing

hero (hero):

That's the hard way to do it. And geez, @satellite you're a genius at typing so fast.

OpenStudy (anonymous):

copy and past is all i did

OpenStudy (anonymous):

and if you have an easier way i am all eyes

OpenStudy (anonymous):

*paste

hero (hero):

You simply reduce one of the expressions to a mixed number. That way it only has one x value. If you wanted to, you could reduce both of them that way. For example: \[\frac{x - 7}{x + 3} = \frac{x + 3 - 10}{x + 3} = \frac{x + 3}{x + 3} - \frac{10}{x + 3} = 1 - \frac{10}{x + 3}\]

hero (hero):

Now you have one x to insert g(x) into f(x)

hero (hero):

\[f(g(x)) = 1 - \frac{10}{\frac{-3x - 7}{x -1} + 3}\]

OpenStudy (anonymous):

i can't wait to see this ...

hero (hero):

It's easier to do on paper than type it out. I've already done this before. I helped another student with the exact same kind of problem.

hero (hero):

But I at least was able to help reduce it so that you're inserting g(x) only once.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

there is no such thing as a free lunch

OpenStudy (anonymous):

ill try to solve it

OpenStudy (anonymous):

\[=\frac{\frac{-3x-7}{x-1}-7}{\frac{-3x-7}{x-1}+3}\times \frac{x-1}{x-1}\] \[=\frac{-3x-7-7(x-1)}{-3x-7+3(x-1)}\]

OpenStudy (anonymous):

distribute?

OpenStudy (anonymous):

is there a typo here?

OpenStudy (anonymous):

oh no it is right yeah distribute combine like terms, etc you will get it but i do like @hero method

hero (hero):

Basically, I re-wrote it like this: \[1 - 10 \div \left(\frac{-3x - 7}{x - 1} + 3\right)\]\[1 - 10 \div \left(\frac{-3x - 7 + 3(x -1)}{x-1}\right)\]\[1 - 10 \times \left(\frac{x - 1}{-10}\right)\]\[1 + x - 1\]

OpenStudy (anonymous):

but you are supposed to get \(x\)

hero (hero):

\[1 - 1 + x = x\]

OpenStudy (anonymous):

how do you get the 10

OpenStudy (anonymous):

i am just funnin

OpenStudy (anonymous):

the 10 was a trick

hero (hero):

@melchar, look at my second post.

OpenStudy (anonymous):

thank you!